# Natural Ultralight Dark Matter: The Quadratic Twin

### v1

@@ I Introduction
The dark matter (DM) problem has been a major driver of research in particle physics, cosmology, and more recently, in atomic, molecular, and optical physics. A large part of this endeavor focuses on the exciting possibility that DM consists of an ultralight (or sub-eV) degree of freedom, whose coherent oscillations account for the observed DM density [1, 2]. This scenario has far-reaching experimental implications, with search strategies ranging from galaxy dynamics to tabletop experiments. For a recent review see [3].
@@
To illustrate this point, consider an ULDM field $\phi$ with a linear or a quadratic coupling to the electron mass operator, denoted $d_{m_{e}}^{(1)}/M_{\text{Pl}}$ and $d_{m_{e}}^{(2)}/(2M_{\text{Pl}}^{2})$, respectively, where $M_{\text{Pl}}$ is the reduced Planck mass. Setting for instance $m_{\phi}=10^{-15}\,$eV, existing experimental bounds on these couplings are $d_{m_{e}}^{(1)}\lesssim 10^{-3}$ and $d_{m_{e}}^{(2)}\lesssim 10^{12}$, respectively, see e.g.  [5]. For the linear case, current experiments probe cutoffs of $\mathcal{O}(100)\,$GeV, which would imply new physics at the weak scale that has not been discovered. For quadratically coupled ULDM, the situation is even worse since currently the best experiments are sensitive to SM cutoffs of $\mathcal{O}(10)\,$eV.
@@
Such fine-tuning problems are common to many scalar and pseudoscalar ULDM theories [19, 20, 21, 22, 23, 24, 25, 26, 27, 28], and have inspired model-building solutions for various classes of models. ULDM theories where the DM field is a pseudo-Nambu-Goldstone boson (pNGB) are often considered in the literature because the pNGB mass is protected by an approximate shift-symmetry. A prominent model is the QCD axion [29, 30, 31], which can play the role of the DM and also solve the strong CP problem. But even with the protection of the approximate shift-symmetry, most experiments are typically not sensitive enough to probe its natural parameter space, and hence require further tuning. The only natural model [32, 33] (see also [34]) populating the region currently probed by experiments as they approach the QCD axion line, envokes $N$ copies of the SM that are related by a discrete $\mathbb{Z}_{N}$ symmetry, which conspire to protect the QCD axion mass. The same mechanism has also been used for natural ULDM with linear couplings to the SM in [35]. The dilaton has also been proposed as an ULDM candidate [36, 37, 38, 39], but due to its generically large quartic, its early universe behavior is in tension with observations in standard theoretical setups [40]. While a consistent theory of the dilaton ULDM model was proposed in [41], its natural parameter space is also well beyond the reach of any experiment.
@@
This work aims to alleviate the naturalness problem of ULDM theories with quadratic couplings large enough to be detected by ongoing and future experiments. Inspired by the twin-Higgs mechanism [42], we consider theories where the ULDM fields are pNGBs whose masses are protected by a mirror-$\mathbb{Z}_{2}$ symmetry relating the SM sector to a twin sector, which is a copy of the SM. As shown in Fig. 1, this theoretical construct opens up roughly 25 orders of magnitude in the parameter space of couplings for DM masses $m_{\phi}\sim(10^{-20}-10^{-15})\,\text{eV}$, which puts natural models at the doorstep of ongoing and near-future experiments.
@@ II The quadratic twin mechanism
[…]
@@
[…]
While the degree of divergence remains the same [^2] This is unlike the twin-Higgs mechanism [42], where the explicit breaking arises from renormalizable interactions, and the mass correction is $\log$-divergent instead. compared to the case without twin contributions in Eq. (4), the mass correction is now quadratic in the coupling. Since the pNGB is coupled extremely weakly to the SM, this correction is typically many orders of magnitude smaller than the contribution in Eq. (4).
@@
[…]
$F_{\mu\nu}$, $G_{a,\mu\nu}$ are the photon and gluon field strengths respectively, $\beta_{g}$ is the beta-function of the strong coupling constant $g$, $e$ the electromagnetic coupling constant, $\gamma_{m_{i}}$ the anomalous dimension and $\psi_{i}$ are the SM fermions. We parameterize the dimensionless couplings $d_{e}^{(2)},d_{g}^{(2)},d_{m_{i}}^{(2)}$ following the convention of [38]. The suppression of the pNGB mass correction in our model, relative to the SM contribution is therefore
@@
Finally, we compare the quadratic twin ULDM to theories of QCD axions [29, 30, 31]. The couplings of the quadratic twin ULDM to the SM and twin sector explicitly break the symmetry group $G$, and are not necessarily linked to any chiral anomaly. In contrast, the QCD axion necessarily breaks the PQ symmetry by the anomaly, which induces similar quadratic couplings to SM particles and the axion mass (see e.g. [10, 50, 51, 52, 53]). Notably, the QCD axion also has derivative couplings from either the anomaly (as in KSVZ models [54, 55]) or the UV (as in DFSZ models [56, 57]). For the quadratic twin, no derivative couplings are generated if the explicit breaking is not related to a chiral anomaly.
III $U(1)$ model
@@ IV Non-abelian groups
Our model can be straightforwardly extended to non-abelian groups. [^3] Interestingly, a similar scenario was recently considered in [58] and shown to potentially give rise to oscillating exotic meson decays. Here we focus on the complex field $\Sigma$ in the fundamental of $G=U(2)$. The $U(2)$ is spontaneously broken to a $U(1)$ subgroup by the VEV $\left\langle|\Sigma|^{2}\right\rangle=f^{2}$ and explicitly broken to $H_{\text{SM}}\times H_{\text{twin}}=U(1)_{\text{SM}}\times U(1)_{\text{twin}}$ by couplings to the SM and twin sectors. This symmetry-breaking pattern leads to 3 pNGBs, which we parameterize as
@@
Due to the mirror-$\mathbb{Z}_{2}$ symmetry, the mass correction to ϕ±superscriptitalic-$\phi^{\pm}$ at linear order in the coupling also exactly cancels in this case. The leading order again arises at quadratic order in the coupling and gives, as for the abelian model, a mass contribution of
@@ V Phenomenology
where $d_{m_{e}}^{(2)}$ can be positive or negative. This mass shift within the Earth (or any other compact object with $n_{e}\neq 0$) leads to screening of the DM field amplitude while gradients thereof are generically enhanced. This has been studied extensively for light QCD axions [4, 7, 8]. We show the details of the calculations for both signs of the coupling in App. B, taking into account finite momentum boundary conditions. The bounds sensitive to amplitudes, such as from clocks and interferometers, can be affected by screening in parts of the parameter space. In Fig. 1, we show the expected bound with solid lines while keeping the Goldstone naturalness bound as dashed lines of the same color. Similarly, in parts of the parameter space, experiments sensitive to DM field gradients, such as EP tests, are affected. In Fig. 1 we show the bounds from MICROSCOPE for positive and negative $d_{m_{e}}^{(2)}$. The resonance features are characteristic of negatively coupled quadratic scalars, assuming a perfectly spherical Earth with homogeneous density [7, 8], see App. B for further details.
The natural region of quadratic twin ULDM can be extensively searched for in the future with ${{}^{229}\rm Th}$ nuclear clocks [50, 18]. Assuming a relative sensitivity of $\delta\alpha/\alpha\sim 10^{-24}$ and $d_{e}\sim(\alpha/4\pi)\times d_{m_{e}}^{(2)}$, we show the nuclear clock projection in Fig. 1 as gray dashed. Stochasticity, not taken into account here, may enhance the reach even further, see e.g. [51].
In the case of negative coupling, the mass can become tachyonic inside finite-density objects, which can source the scalar field whether or not it constitutes any fraction of the DM. This has been extensively studied in the context of (light) QCD axions, where it can lead to extremely stringent constraints [9, 10, 59, 11, 12, 13, 14]. For instance, [11] showed that the presence of a quadratically coupled light QCD axion can change the mass-radius relationship of white dwarfs which is inconsistent with observations. These constraints arise when the change of the fermion mass due to the sourced field is relatively large, as is the case for QCD axions, and disappear for a coupling to electrons when $\delta m_{e}/m_{e}<10^{-3}$ [11, 60, 61].
@@
Similarly, quadratic twin ULDM also avoids bounds from Big Bang nucleosynthesis [62], since they rely on relatively large changes to the electron mass.
@@ VI Conclusions and Discussion
We have presented a model of naturally ultralight, quadratically coupled scalar dark matter. Inspired by the twin-Higgs mechanism [42], we considered a pNGB of some spontaneously broken global symmetry $G$. We introduced couplings to our SM and a twin sector, related by a mirror-$\mathbb{Z}_{2}$ symmetry, which explicitly breaks $G$ to $H_{\text{SM}}\times H_{\text{twin}}$.
@@
Finally, we want to comment on the cosmological history of quadratic twin ULDM. While it can, in principle, be produced by misalignment, the quadratic twin ULDM comes with the usual difficulties of the twin-Higgs cosmology, see e.g. [63, 64]. The mirror-$\mathbb{Z}_{2}$ has to be broken in the early universe since constraints on the effective number of degrees of freedom do not allow reheating the twin SM. Therefore, the thermal mass induced by the SM bath is not protected by the mirror-$\mathbb{Z}_{2}$, which can lead to either larger, smaller, or even tachyonic masses during the early Universe evolution, depending on the sign of the coupling. The resulting cosmological history can have rich phenomenology, with different regimes for ULDM equation of state and potential phase transitions. These consideration might be important for possible production mechanism of quadratic twin ULDM.
@@ Appendix A Goldstone naturalness
$\delta\lambda\simeq\frac{d^{(2)}_{m_{\psi}}m_{\psi}^{2}\Lambda^{2}}{64\pi^{2}M_{\text{Pl}}^{4}}\lesssim\frac{\delta m_{\phi}^{2}}{f^{2}}\implies f\lesssim\sqrt{\frac{2}{d^{(2)}_{m_{\psi}}}}M_{\text{Pl}}.$ (20)
@@ Appendix B Screening and sourcing
In this appendix, we discuss the phenomenological implications of a density-dependent scalar mass. We first consider screening, which suppresses the DM amplitude but enhances gradients, as shown in [4, 7, 8]. Second, we discuss the scenario in which the scalar mass turns tachyonic which leads to sourcing [9, 10, 65, 66, 11, 12, 13].
@@
We closely follow [7] and define momentum inside the Earth as $q$ and the momentum outside Earth as $k$:
$k^{2}\equiv\omega^{2}-m_{\phi}^{2}$ (inside) (21a)
$q^{2}\equiv\omega^{2}-m_{\phi}^{2}-\delta m_{\phi}^{2}$ (outside) (21b)
@@
Let us recap where in the parameter space we expect these effects to be important, see [7]. As can be seen in the upper left panel of Fig. 2, the mass change needs to be significant (for both signs of the coupling) in order for an appreciable effect to occur. Two conditions must be satisfied for the effect to occur which are $qR\gtrsim\mathcal{O}(1)$ and $|q^{2}|>|k^{2}|$, which can both be approximated as $\delta m_{\phi}R>1$ and $|(\delta m_{\phi}R)^{2}|>(kR)^{2}$. These can be roughly interpreted as requiring the object to be large and dense enough, respectively.
Additionally, there is a region where gradients are enhanced. This can be estimated in the limit where $kR<1$, where the gradient scales as $\partial_{r}\phi/(k\phi_{0})\sim(qR)^{2}/(3kR)\sim(\delta m_{\phi}R)^{2}/(3kR)>1$. We show these regions for positive (left panel) and negative (right panel) of $d_{m_{e}}^{(2)}$ in Fig. 3: Above the blue line marks the region in which the amplitude is suppressed, while the region within the green line marks the region where the gradient is enhanced. The dark purple line corresponds to the region where the scalar mass is tachyonic, and the solutions presented in [7] are no longer valid. This is because the field is displaced from its in-vacuum minimum and stabilized at $\phi_{n}=\pi f/2$, see Fig. 5. We leave the evaluation of the DM profile in this case for future work, see also discussion below.
@@
Figure 3: Same as Fig. 1 but including regions where for both signs ($d_{m_{e}}^{(2)}>0$ in the left panel and $d_{m_{e}}^{(2)}<0$ in the right panel) the DM field amplitude is screened (above blue line), DM field gradients are enhanced (above green line) and where the quadratic twin ULDM is sourced within Earth (region above the dark purple line in the right panel for $d_{m_{e}}^{(2)}<0$).
@@
$\left(\frac{\delta m_{e}}{m_{e}}\right)_{\text{lin}}$ $=\frac{d_{m_{e}}^{(1)}}{M_{\text{Pl}}}\phi_{0}\cos\left(m_{\phi}t\right)=\frac{d_{m_{e}}^{(1)}}{M_{\text{Pl}}}\frac{\sqrt{2\rho_{\text{DM}}}}{m_{\phi}}\cos\left(m_{\phi}t\right),$ (25)
@@
Therefore, to find the bound on the quadratic coupling one can take the linear result with $m_{\phi}\to m_{\phi}/2$ and
$d_{m_{e}}^{(1)}\to d_{m_{e}}^{(2)}\frac{M_{\text{Pl}}}{\sqrt{2\rho_{\text{DM}}}}.$ (26)
@@
which enters the effective quadratic coupling. The bound obtained without screening is shown in Fig. 1 and Fig. 3 in dashed lines, while the screening-accounted bound is shown in solid lines. Notice that we focus on the regime where $kR<1$, which corresponds to $m_{\phi}\lesssim 10^{-11}\,\text{eV}$, since reliable calculations can only be done in this regime [7]. Therefore, we are not showing bounds from [67] and [68] which lie in the $m_{\phi}>10^{-11}\,\text{eV}$ regime.
Contrary to the reduced sensitivity of amplitude-dependent experiments, gradient-dependent experiments would generically see an enhancement $\sim 1/(kR)^{2}$. This fact has recently been exploited for EP tests such as the MICROSCOPE experiment which is sensitive to $\sim\phi\nabla\phi$ in [49] for the axion coupling, using the profiles calculated in [7].
@@
where $\phi_{\text{out}}$ is a function of $m_{\phi}R_{\oplus},m_{\phi}R_{\text{mic}}$ and $m_{\phi}(R_{\text{mic}}-R_{\oplus})$ [7] and $Q_{i}^{M}$ are the dilatonic charges. For the electron coupling, the dilatonic charges for $d_{m_{e}}^{(2)}$ are given by [38]
@@
In the case that $kR_{\oplus}<1$ it is sufficient to only consider the $\ell=0$ mode. The bound from Microscope is roughly $\eta>10^{-14}$ and plot the Eötvös parameter as a function of $d_{m_{e}}^{(2)}$ in Fig. 4.
@@
At these densities the point $\phi_{0}=0$ ceases to be a minimum and new minima appear at $\phi_{0}^{n}=\pm\pi f/2$. In Fig. 5 we show the non-relativistic 1-loop Coleman-Weinberg potential for $n_{e}=0$ (blue) as well as sub- (orange) and super-critical densities (green). As we can see, the minimum at $\phi=0$ gets destabilized and the minima at ±$\pm\pi f/2$ are energetically preferred.
@@
Note that no linear coupling is generated in regions of the object where the field sits in the finite density minimum of the potential. If gradient effects are important the scalar can be between the two minima within regions of the star and linear couplings would be induced. These would give rise to strong bounds from e.g. stellar cooling [60], which we do not consider here.
Even without linear couplings, sourcing can lead to large observable effects. In [11, 60, 61] it is shown that strong bounds can be derived if the mass reduction of the electron is large enough. Concretely, that implies that the bound from mass-radius observations of White Dwarfs disappears for mass changes smaller than $\delta m_{e}/m_{e}<10^{-3}$. In our case, if we identify the field range with the cutoff $f=\Lambda=10\,\text{TeV}$, we find that the electron mass change is tiny except for the largest values of $d_{m_{e}}^{(2)}$, and therefore we do not show bounds from modified white dwarf mass-radius curves.

### v2

@@ I Introduction
The dark matter (DM) problem has been a major driver of research in particle physics, cosmology, and more recently, in atomic, molecular, and optical physics. A large part of this endeavor focuses on the exciting possibility that DM consists of an ultralight (or sub-eV) degree of freedom, whose coherent oscillations account for the observed DM density [Preskill:1982cy, Dine:1982ah]. This scenario has far-reaching experimental implications, with search strategies ranging from galaxy dynamics to tabletop experiments. For a recent review see [Antypas:2022asj].
@@
To illustrate this point, consider an ULDM field $\phi$ with a linear or a quadratic coupling to the electron mass operator, denoted $d_{m_{e}}^{(1)}/M_{\text{Pl}}$ and $d_{m_{e}}^{(2)}/(2M_{\text{Pl}}^{2})$, respectively, where $M_{\text{Pl}}$ is the reduced Planck mass. Setting for instance $m_{\phi}=10^{-15}\,$eV, existing experimental bounds on these couplings are $d_{m_{e}}^{(1)}\lesssim 10^{-3}$ and $d_{m_{e}}^{(2)}\lesssim 10^{12}$, respectively, see e.g.  [Banerjee:2022sqg]. For the linear case, current experiments probe cutoffs of $\mathcal{O}(100)\,$GeV, which would imply new physics at the weak scale that has not been discovered. For quadratically coupled ULDM, the situation is even worse since currently the best experiments are sensitive to SM cutoffs of $\mathcal{O}(10)\,$eV.
@@
Such fine-tuning problems are common to many scalar and pseudoscalar ULDM theories [Froggatt:1978nt, Leurer:1992wg, Leurer:1993gy, Nelson:1983zb, Nelson:1984hg, Barr:1984qx, Piazza:2010ye, Graham:2015cka, Calibbi:2016hwq, Dine:2024bxv], and have inspired model-building solutions for various classes of models. ULDM theories where the DM field is a pseudo-Nambu-Goldstone boson (pNGB) are often considered in the literature because the pNGB mass is protected by an approximate shift-symmetry. A prominent model is the QCD axion [Peccei:1977hh, Wilczek:1977pj, Weinberg:1977ma], which can play the role of the DM and also solve the strong CP problem. But even with the protection of the approximate shift-symmetry, most experiments are typically not sensitive enough to probe its natural parameter space, and hence require further tuning. The only natural model [Hook_2018b, Banerjee:2025kov] (see also [DiLuzio:2021pxd]) populating the region currently probed by experiments as they approach the QCD axion line, envokes $N$ copies of the SM that are related by a discrete $\mathbb{Z}_{N}$ symmetry, which conspire to protect the QCD axion mass. The same mechanism has also been used for natural ULDM with linear couplings to the SM in [Brzeminski:2020uhm]. The dilaton has also been proposed as an ULDM candidate [Cho:1998js, Damour:2010rm, Damour:2010rp, Arvanitaki:2014faa], but due to its generically large quartic, its early universe behavior is in tension with observations in standard theoretical setups [Hubisz:2024hyz]. While a consistent theory of the dilaton ULDM model was proposed in [Banerjee:2025uwn], its natural parameter space is also well beyond the reach of any experiment.
@@
This work aims to alleviate the naturalness problem of ULDM theories with quadratic couplings large enough to be detected by ongoing and future experiments. Inspired by the twin-Higgs mechanism [Chacko:2005pe], we consider theories where the ULDM fields are pNGBs whose masses are protected by a mirror-$\mathbb{Z}_{2}$ symmetry relating the SM sector to a twin sector, which is a copy of the SM. As shown in Fig. 1, this theoretical construct opens up roughly 25 orders of magnitude in the parameter space of couplings for DM masses $m_{\phi}\sim(10^{-20}-10^{-15})\,\text{eV}$, which puts natural models at the doorstep of ongoing and near-future experiments.
@@ II The quadratic twin mechanism
$\text{SM:}\quad\delta m_{\text{pNGB}}^{2}\simeq\frac{d_{m_{\psi}}^{(2)}m_{\psi}^{2}\Lambda^{2}}{16\pi^{2}M_{\text{Pl}}^{2}},$ (4)
@@
$\text{SM+twin:}\quad\delta m_{\text{pNGB}}^{2}=\frac{(d_{m_{\psi}}^{(2)})^{2}m_{\psi}^{2}f^{2}}{16\pi^{2}M_{\text{Pl}}^{4}}\Lambda^{2}.$ (7)
While the degree of divergence remains the same [^2] This is unlike the twin-Higgs mechanism [Chacko:2005pe], where the explicit breaking arises from renormalizable interactions, and the mass correction is $\log$-divergent instead. compared to the case without twin contributions in Eq. (4), the mass correction is now quadratic in the coupling. Since the pNGB is coupled extremely weakly to the SM, this correction is typically many orders of magnitude smaller than the contribution in Eq. (4).
@@
$\bigg.-\left(d_{m_{i}}^{(2)}+\gamma_{m_{i}}d_{g}^{(2)}\right)m_{i}\bar{\psi}_{i}\psi_{i}+\text{h.c.}\bigg]+\text{twin-sector}.$
$F_{\mu\nu}$, $G_{a,\mu\nu}$ are the photon and gluon field strengths respectively, $\beta_{g}$ is the beta-function of the strong coupling constant $g$, $e$ the electromagnetic coupling constant, $\gamma_{m_{i}}$ the anomalous dimension and $\psi_{i}$ are the SM fermions. We parameterize the dimensionless couplings $d_{e}^{(2)},d_{g}^{(2)},d_{m_{i}}^{(2)}$ following the convention of [Damour:2010rp]. The suppression of the pNGB mass correction in our model, relative to the SM contribution is therefore
@@
Finally, we compare the quadratic twin ULDM to theories of QCD axions [Peccei:1977hh, Wilczek:1977pj, Weinberg:1977ma]. The couplings of the quadratic twin ULDM to the SM and twin sector explicitly break the symmetry group $G$, and are not necessarily linked to any chiral anomaly. In contrast, the QCD axion necessarily breaks the PQ symmetry by the anomaly, which induces similar quadratic couplings to SM particles and the axion mass (see e.g. [Balkin:2020dsr, Kim:2022ype, Kim:2023pvt, Beadle:2023flm, Springmann:2024mjp]). Notably, the QCD axion also has derivative couplings from either the anomaly (as in KSVZ models [Kim:1979if, Shifman:1979if]) or the UV (as in DFSZ models [Dine:1981rt, Zhitnitsky:1980tq]). For the quadratic twin, no derivative couplings are generated if the explicit breaking is not related to a chiral anomaly.
[…]
@@ IV Non-abelian groups
Our model can be straightforwardly extended to non-abelian groups. [^3] Interestingly, a similar scenario was recently considered in [Bigaran:2025uzn] and shown to potentially give rise to oscillating exotic meson decays. Here we focus on the complex field $\Sigma$ in the fundamental of $G=U(2)$. The $U(2)$ is spontaneously broken to a $U(1)$ subgroup by the VEV $\left\langle|\Sigma|^{2}\right\rangle=f^{2}$ and explicitly broken to $H_{\text{SM}}\times H_{\text{twin}}=U(1)_{\text{SM}}\times U(1)_{\text{twin}}$ by couplings to the SM and twin sectors. This symmetry-breaking pattern leads to 3 pNGBs, which we parameterize as
@@
Due to the mirror-$\mathbb{Z}_{2}$ symmetry, the mass correction to $\phi^{\pm}$ at linear order in the coupling also exactly cancels in this case. The leading order again arises at quadratic order in the coupling and gives, as for the abelian model, a mass contribution of
@@ V Phenomenology
where $d_{m_{e}}^{(2)}$ can be positive or negative. This mass shift within the Earth (or any other compact object with $n_{e}\neq 0$) leads to screening of the DM field amplitude while gradients thereof are generically enhanced. This has been studied extensively for light QCD axions [Hees:2018fpg, Banerjee:2025dlo, delCastillo:2025rbr]. We show the details of the calculations for both signs of the coupling in App. B, taking into account finite momentum boundary conditions. The bounds sensitive to amplitudes, such as from clocks and interferometers, can be affected by screening in parts of the parameter space. In Fig. 1, we show the expected bound with solid lines while keeping the Goldstone naturalness bound as dashed lines of the same color. Similarly, in parts of the parameter space, experiments sensitive to DM field gradients, such as EP tests, are affected. In Fig. 1 we show the bounds from MICROSCOPE for positive and negative $d_{m_{e}}^{(2)}$. The resonance features are characteristic of negatively coupled quadratic scalars, assuming a perfectly spherical Earth with homogeneous density [Banerjee:2025dlo, delCastillo:2025rbr], see App. B for further details.
The natural region of quadratic twin ULDM can be extensively searched for in the future with $^{229}\rm Th$} nuclear clocks [Kim:2022ype, Caputo:2024doz]. Assuming a relative sensitivity of $\delta\alpha/\alpha\sim 10^{-24}$ and $d_{e}\sim(\alpha/4\pi)\times d_{m_{e}}^{(2)}$, we show the nuclear clock projection in Fig. 1 as gray dashed. Stochasticity, not taken into account here, may enhance the reach even further, see e.g. [Kim:2023pvt].
In the case of negative coupling, the mass can become tachyonic inside finite-density objects, which can source the scalar field whether or not it constitutes any fraction of the DM. This has been extensively studied in the context of (light) QCD axions, where it can lead to extremely stringent constraints [Hook:2017psm, Balkin:2020dsr, Zhang:2021mks, Balkin:2022qer, Balkin:2023xtr, Gomez-Banon:2024oux, Kumamoto:2024wjd]. For instance, [Balkin:2022qer] showed that the presence of a quadratically coupled light QCD axion can change the mass-radius relationship of white dwarfs which is inconsistent with observations. These constraints arise when the change of the fermion mass due to the sourced field is relatively large, as is the case for QCD axions, and disappear for a coupling to electrons when $\delta m_{e}/m_{e}<10^{-3}$ [Balkin:2022qer, Bartnick:2025lbv, Bartnick:2025lbg].
@@
Similarly, quadratic twin ULDM also avoids bounds from Big Bang nucleosynthesis [Bouley:2022eer], since they rely on relatively large changes to the electron mass.
@@ VI Conclusions and Discussion
We have presented a model of naturally ultralight, quadratically coupled scalar dark matter. Inspired by the twin-Higgs mechanism [Chacko:2005pe], we considered a pNGB of some spontaneously broken global symmetry $G$. We introduced couplings to our SM and a twin sector, related by a mirror-$\mathbb{Z}_{2}$ symmetry, which explicitly breaks $G$ to $H_{\text{SM}}\times H_{\text{twin}}$.
@@
Finally, we want to comment on the cosmological history of quadratic twin ULDM. While it can, in principle, be produced by misalignment, the quadratic twin ULDM comes with the usual difficulties of the twin-Higgs cosmology, see e.g. [Craig:2016lyx, Chacko:2018vss]. The mirror-$\mathbb{Z}_{2}$ has to be broken in the early universe since constraints on the effective number of degrees of freedom do not allow reheating the twin SM. Therefore, the thermal mass induced by the SM bath is not protected by the mirror-$\mathbb{Z}_{2}$, which can lead to either larger, smaller, or even tachyonic masses during the early Universe evolution, depending on the sign of the coupling. The resulting cosmological history can have rich phenomenology, with different regimes for ULDM equation of state and potential phase transitions. These consideration might be important for possible production mechanism of quadratic twin ULDM.
@@ Appendix A Goldstone naturalness
$\delta\lambda\simeq\frac{(d^{(2)}_{m_{\psi}})^{2}m_{\psi}^{2}\Lambda^{2}}{64\pi^{2}M_{\text{Pl}}^{4}}\lesssim\frac{\delta m_{\phi}^{2}}{f^{2}}\implies f\lesssim\sqrt{\frac{2}{d^{(2)}_{m_{\psi}}}}M_{\text{Pl}}.$ (20)
@@ Appendix B Screening and sourcing
In this appendix, we discuss the phenomenological implications of a density-dependent scalar mass. We first consider screening, which suppresses the DM amplitude but enhances gradients, as shown in [Hees:2018fpg, Banerjee:2025dlo, delCastillo:2025rbr]. Second, we discuss the scenario in which the scalar mass turns tachyonic which leads to sourcing [Hook:2017psm, Balkin:2020dsr, Balkin:2021wea, Balkin:2021zfd, Balkin:2022qer, Balkin:2023xtr, Gomez-Banon:2024oux].
@@
We closely follow [Banerjee:2025dlo] and define momentum inside the Earth as $q$ and the momentum outside Earth as $k$:
$k^{2}\equiv\omega^{2}-m_{\phi}^{2}$ (outside) (21a)
$q^{2}\equiv\omega^{2}-m_{\phi}^{2}-\delta m_{\phi}^{2}$ (inside) (21b)
@@
Let us recap where in the parameter space we expect these effects to be important, see [Banerjee:2025dlo]. As can be seen in the upper left panel of Fig. 2, the mass change needs to be significant (for both signs of the coupling) in order for an appreciable effect to occur. Two conditions must be satisfied for the effect to occur which are $qR\gtrsim\mathcal{O}(1)$ and $|q^{2}|>|k^{2}|$, which can both be approximated as $\delta m_{\phi}R>1$ and $|(\delta m_{\phi}R)^{2}|>(kR)^{2}$, with $\delta m_{\phi}$ the absolute magnitude of the density induced mass. These can be roughly interpreted as requiring the object to be large and dense enough, respectively.
Additionally, there is a region where gradients are enhanced. This can be estimated in the limit where $kR<1$, where the gradient scales as $\partial_{r}\phi/(k\phi_{0})\sim(qR)^{2}/(3kR)\sim(\delta m_{\phi}R)^{2}/(3kR)>1$. We show these regions for positive (left panel) and negative (right panel) of $d_{m_{e}}^{(2)}$ in Fig. 3: Above the blue line marks the region in which the amplitude is suppressed, while the region within the green line marks the region where the gradient is enhanced. The dark purple line corresponds to the region where the scalar mass is tachyonic, and the solutions presented in [Banerjee:2025dlo] are no longer valid. This is because the field is displaced from its in-vacuum minimum and stabilized at $\phi_{n}=\pi f/2$, see Fig. 5. We leave the evaluation of the DM profile in this case for future work, see also discussion below.
@@
Figure 3: Same as Fig. 1 but including regions where for both signs ($d_{m_{e}}^{(2)}>0$ in the left panel and $d_{m_{e}}^{(2)}<0$ in the right panel) the DM field amplitude is screened (above blue line), DM field gradients are enhanced (above green line) and where the quadratic twin ULDM is sourced within Earth (region above the dark purple line in the right panel for $d_{m_{e}}^{(2)}<0$). In the right panel the light blue line is the bound from MICROSCOPE ([Banerjee:2025dlo, Gue:2025nxq]), which is not valid within the dark purple region in which the field is sourced. In this region the effective mass is larger than the vacuum mass and positive since one expands around the finite density minimum.
@@
$\left(\frac{\delta m_{e}}{m_{e}}\right)_{\text{lin}}$ $=\frac{d_{m_{e}}^{(1)}}{M_{\text{Pl}}}\phi_{0}\cos(m_{\phi}t)=\frac{d_{m_{e}}^{(1)}}{M_{\text{Pl}}}\frac{\sqrt{2\rho_{\text{DM}}}}{m_{\phi}}\cos(m_{\phi}t),$ (25)
@@
Therefore, to find the bound on the quadratic coupling we can make the replacements
$m_{\phi,\rm quad}$ $=\frac{m_{\phi,\rm lin}}{2},$ (26)
$d_{m_{e}}^{(2)}$ $=d_{m_{e}}^{(1)}\sqrt{\frac{2}{\rho_{\rm DM}}}M_{\text{Pl}}m_{\phi,\rm quad}.$
@@
which enters the effective quadratic coupling. The bound obtained without screening is shown in Fig. 1 and Fig. 3 in dashed lines, while the screening-accounted bound is shown in solid lines. Notice that we focus on the regime where $kR<1$, which corresponds to $m_{\phi}\lesssim 10^{-11}\,\text{eV}$, since reliable calculations can only be done in this regime [Banerjee:2025dlo]. Therefore, we are not showing bounds from [Antypas:2019qji] and [Savalle:2020vgz] which lie in the $m_{\phi}>10^{-11}\,\text{eV}$ regime.
Contrary to the reduced sensitivity of amplitude-dependent experiments, gradient-dependent experiments would generically see an enhancement $\sim 1/(kR)^{2}$. This fact has recently been exploited for EP tests such as the MICROSCOPE experiment which is sensitive to $\sim\phi\nabla\phi$ in [Gue:2025nxq] for the axion coupling, using the profiles calculated in [Banerjee:2025dlo].
@@
where $\phi_{\text{out}}$ is a function of $m_{\phi}R_{\oplus},m_{\phi}R_{\text{mic}}$ and $m_{\phi}(R_{\text{mic}}-R_{\oplus})$ [Banerjee:2025dlo] and $Q_{i}^{M}$ are the dilatonic charges. Here we use $R_{\text{mic}}\simeq 710\,\rm km$. For the electron coupling, the dilatonic charges for $d_{m_{e}}^{(2)}$ are given by [Damour:2010rp]
@@
In the case that $kR_{\oplus}<1$ it is sufficient to only consider the $\ell=0$ mode. The bound from Microscope is roughly $|\eta|<10^{-14}$ and plot the Eötvös parameter as a function of $d_{m_{e}}^{(2)}$ in Fig. 4.
@@
At these densities the point $\phi_{0}=0$ ceases to be a minimum and new minima appear at $\phi_{0}^{n}=\pm\pi f/2$. In Fig. 5 we show the non-relativistic 1-loop Coleman-Weinberg potential for $n_{e}=0$ (blue) as well as sub- (orange) and super-critical densities (green). As we can see, the minimum at $\phi=0$ gets destabilized and the minima at $\pm\pi f/2$ are energetically preferred.
@@
Note that no linear coupling is generated in regions of the object where the field sits in the finite density minimum of the potential. If gradient effects are important the scalar can be between the two minima within regions of the star and linear couplings would be induced. These would give rise to strong bounds from e.g. stellar cooling [Bartnick:2025lbv], which we do not consider here.
Even without linear couplings, sourcing can lead to large observable effects. In [Balkin:2022qer, Bartnick:2025lbv, Bartnick:2025lbg] it is shown that strong bounds can be derived if the mass reduction of the electron is large enough. Concretely, that implies that the bound from mass-radius observations of White Dwarfs disappears for mass changes smaller than $\delta m_{e}/m_{e}<10^{-3}$. In our case, if we identify the field range with the cutoff $f=\Lambda=10\,\text{TeV}$, we find that the electron mass change is tiny except for the largest values of $d_{m_{e}}^{(2)}$, and therefore we do not show bounds from modified white dwarf mass-radius curves.